# Graphite: A GPU-Accelerated Mixed-Precision Graph Optimization Framework

### v1

@@ IV-B Batching
In general, properties for a constraint descriptor include which vertex descriptors are involved, the data types (classes) of the observations and any non-optimizable data, the loss function used, and the differentiation mode (discussed in Section IV-C and Section IV-D). These properties also specify how to compute the residual for a constraint, and how to compute the Jacobians when using analytic differentiation (Section IV-C). Figure 4 shows this for the point optimization problem. The user specifies a residual of size 1 (Line 3), specifies the point vertex descriptor (Line 4), chooses the default loss function (Line 7), and chooses to use automatic differentiation (Line 8), so no Jacobian evaluation function is specified. Additionally, Lines 10-18 tell the descriptor how to compute the residual based on the squared distance between the point and the radius.
@@
We specify properties of descriptors in this way because it lets users declare the behaviour of their types and automatically generate optimized GPU code, without relying on runtime lookups (as observed in some CPU-based optimizers [3]) to decide how each item in the graph is processed.
// Factor traits for the circle constraint
template <typename T, typename S> struct CircleFactorTraits {
static constexpr size_t dimension = 1;
using VertexDescriptors = std::tuple<PointDescriptor<T, S>>;
using Observation = T;
using Data = unsigned char;
using Loss = DefaultLoss<T, dimension>;
using Differentiation = DifferentiationMode::Auto;
[…]
hd_fn static void error(const D *point, const T *obs, D *error,
const std::tuple<Point<T> *> &vertices,
const Data *data) {
auto x = point[0];
auto y = point[1];
auto r = obs[0];
error[0] = x * x + y * y - r * r;
[…]
using CircleFactor = FactorDescriptor<T, S, CircleFactorTraits<T, S>>;
Figure 4: Properties of a constraint checking a point’s distance from the circumference of a circle with radius $r$.
@@
After the user finishes adding descriptors to the graph and populating them with information about each vertex and constraint, they pass the graph to an optimization algorithm, along with a linear solver (discussed further in Section IV-E). The user optimizes the graph for the point optimization example in Figure 5. First, they decide to fix the value of the last point they created (Line 1) and decide to disable the constraint which was created for the third point by changing its level to 1, as in g2o [3] (Line 2). Next, the user configures the iterative linear solver, by selecting an identity matrix for the preconditioner (Line 3) and then instantiating the solver (Line 4) with their desired parameters. They choose 10 iterations for Levenberg-Marquardt (Line 6), and create one stream (Line 7), since each constraint only has one vertex. On Line 8, they set the optimization level to 0, meaning that any constraints with a level higher than 0, such as the constraint on Line 2, will be excluded from the optimization. Finally, they perform the optimization by calling Levenberg-Marquardt in Lines 9-10.
point_desc.set_fixed(num_points - 1, true); // Set the last vertex as fixed
factor_desc.set_active(2, 0x1); // Disable third constraint for point 2
IdentityPreconditioner<FP, SP> preconditioner;
PCGSolver<FP, SP> solver(50, 1e-6, 10.0, &preconditioner);
[…]
StreamPool streams(1); // chose 1 stream for 1 vertex per constraint
constexpr uint8_t optimization_level = 0;
optimizer::levenberg_marquardt<FP, SP>(&graph, &solver, iterations, 1e-6,
optimization_level, streams);
Figure 5: Example of optimizing noisy 2D points.
IV-C Analytic Differentiation
@@
A typical approach on the CPU is to compute each matrix for a constraint sequentially [3, 4]. However, to take advantage of the many GPU threads available, Graphite computes each Jacobian concurrently in separate streams (we describe streams in Section II). Additionally, Graphite allows these Jacobians to be evaluated dynamically [22], at the point of computation (e.g. matrix-vector multiplication), to enable matrix-free solving methods which use less GPU memory.
IV-D Automatic Differentiation
Another method to compute Jacobians is through forward mode automatic differentiation. For example, Ceres Solver computes the gradient of a residual function using dual numbers, which consist of a real component $a$, for the computed value, and an infinitesimal component $b\varepsilon$, where $\varepsilon^{2}=0$, for the corresponding partial derivative [4]. To extend this for multiple inputs, Ceres implements a data structure called a Jet, which replaces the infinitesimal component with a vector [4]. Ren et. al [15] extend this concept to the GPU by proposing JetVectors, which adopt a structure of arrays data layout to enable coalesced memory transactions for better performance. However, this design is intended for optimization across multiple devices and may require more GPU memory than available on a single device [15].
@@ IV-D Automatic Differentiation
IV-E Mixed-Precision Solver
To enable mixed-precision solving, which enables the optimization to use less memory, Graphite supports the customization of floating point precisions for descriptors and the linear system (Eq. 6). Essentially, these two precisions can be customized according to trade-offs between convergence, runtime performance, and memory usage. A typical approach would be to use a higher precision (double or single precision) for the descriptors to maintain numerical stability, and an equal or lower precision (single precision or brain floating point), to reduce the amount of memory needed to store the Jacobian matrices [32]. In contrast, existing solvers support only one precision at a time [15], or only allow casting the linear system to single precision [4]. As shown in Section V, using lower precisions greatly reduces the overall optimization time and amount of GPU memory required.
As described in Section II, each iteration of Levenberg-Marquardt optimization computes a parameter step $\Delta x$. To compute this quickly on the GPU, we apply preconditioned conjugate gradients (PCG) on the Hessian, with implicit evaluation using the Jacobian matrices [22], since this does not require storing the Hessian, and comprises of algebraic operations which are largely amenable to parallelization. However, directly using PCG with lower precision types can result in poor convergence. To mitigate this, we adopt techniques which are known to help with solving larger systems. These include, clamping values along the Hessian diagonal [4], rescaling the column of each Jacobian block by the square-root inverse of the corresponding Hessian diagonal value [4, 32], and normalizing the PCG residual in each iteration [33].
In applying PCG, to improve convergence, we also use a block Jacobi preconditioner [34], because it is relatively simple to compute and does not require much memory due to its sparse block-diagonal matrix structure. We have observed that storing this preconditioner at too low of a precision can also result in convergence issues. Thus, when a 16-bit precision type such as bfloat16 is specified for the linear system, we compute and store this preconditioner at single or double precision, according to the precision of the graph.
Additionally, since the PCG solver does not require storing the Hessian matrix, we avoid explicitly constructing sparse matrices for the optimization problem. Instead, when performing an operation such as matrix-vector multiplication, the kernel uses the structure of the graph itself. This eliminates additional memory and runtime costs associated with converting and storing matrices in different formats, although each kernel must now be aware of the graph structure and whether a variable is fixed or constraint is inactive.
@@ V Evaluation
Machine Specs
12-core Intel Core i7-12700K CPU @ 4.9 GHz
Desktop 10496-core NVIDIA RTX 3090 GPU @ 1.9 GHz
64 GB RAM
6-core ARM Cortex-A78AE CPU @ 1.7 GHz
Orin Nano 1024-core NVIDIA Ampere GPU @ 1.1 GHz
8 GB RAM
TABLE I: Evaluation machine specifications.
Problem Metric FP64 FP32 FP32-BF16 Ceres-CGNR MegBA-FP64 MegBA-FP32
MSE 0.43 0.43 0.43 0.43 0.43 0.43
Dubrovnik-16 Time 0.22 0.11 0.12 1.64 0.17 0.16
GPU Memory 298 282 278 336 406 352
MSE 1.67 1.67 1.68 1.67 1.67 1.67
Trafalgar-21 Time 0.11 0.069 0.068 0.82 0.098 0.088
GPU Memory 280 274 270 318 326 298
MSE 0.85 0.85 0.85 0.84 0.84 0.84
Ladybug-49 Time 0.063 0.041 0.046 0.66 0.086 0.074
GPU Memory 278 274 270 316 326 300
MSE 0.67 0.67 0.67 0.67 0.66 0.67
Venice-1778 Time 6.00 2.80 2.87 43.71 6.96 6.15
GPU Memory 1944 1190 950 2160 7810 4528
MSE 1.13 1.13 1.14 1.14 1.12 1.13
Final-4585 Time 9.62 4.69 3.75 73.49 13.19 9.82
GPU Memory 3102 1822 1386 3546 13882 7960
MSE 1.50 1.50 1.50 1.51 OOM OOM
Final-13682 Time 26.60 11.65 11.67 210.64 OOM OOM
GPU Memory 9298 5212 3828 10640 OOM OOM
TABLE II: Time (seconds), mean squared error, and GPU memory usage (MiB) for BAL experiments using different precisions (desktop).
@@ 8 GB RAM
Final-13682 Time 26.57 417 38.81
(FP64) GPU Memory 9298 3988 9298
@@
Final-13682 Time 10.97 25.22 11.46
(FP32) GPU Memory 5212 2556 5212
@@
Final-13682 Time 10.64 24.48 11.27
(FP32-BF16) GPU Memory 3828 2500 3828
TABLE III: Time (seconds), mean squared error, and GPU memory usage (MiB) for across differentiation modes (desktop).
Sequence Initial $\chi^{2}$ Graph Size Metric g2o-LDLT g2o-PCG FP64 FP32 FP32-BF16
@@
(Orin Nano) 174070 Co GPU Mem. 4674, 282 4473, 81 4449, 57
TABLE IV: Results of full-inertial bundle adjustment on generated maps. Time given in seconds and GPU memory in MiB.
We evaluate Graphite in terms of the error reduction, runtime performance, and GPU memory usage. First, we compare the solver in different precision modes, using problems from Bundle Adjustment in the Large (BAL) [34], a popular dataset for least squares and bundle adjustment solvers. Next, we compare analytic, dynamic, and automatic differentiation modes. Lastly, we run full-inertial bundle adjustment with Graphite inside ORB-SLAM3 using maps generated from outdoors sequences in TUM-VI, a well-known dataset for evaluating visual-inertial SLAM systems. We run these experiments using a desktop machine equipped with a discrete graphics card, and an NVIDIA Jetson Orin Nano Developer Kit, an embedded robotics platform, to assess the performance of Graphite in high-performance and resource-constrained environments (specifications in Table I).
@@ V-A Mixed-Precision Modes
We evaluate Graphite on BAL datasets using double precision (FP64), single precision (FP32), and single precision mixed with bfloat16 (FP32-BF16) on the desktop machine. We use bfloat16 rather than 16-bit half precision, because it can approximately represent the same range as single precision. For comparison, we also include two GPU-accelerated solvers, Ceres Solver [4], a versatile solver commonly used for SLAM, configured for CUDA CGNR (conjugate gradients on the normal equations), which like our solver, also uses implicit Hessian evaluation, and MegBA [15], a high-performance solver for bundle adjustment with support for single and double precision solving. Where available, we use analytic Jacobians, which have the least runtime overhead. We generate analytic Jacobians for our solver using wrenfold [31] for symbolic differentiation. To allow each optimizer sufficient opportunity to converge, we use a maximum of 50 iterations. For PCG in MegBA, the maximum number of iterations is set to 100, while 10 is used for the other configurations, since they solve the full system rather than a reduced one, which involves more work per iteration. As shown in Table II, the solver is able to use substantially less memory when using BF16 to store the Jacobian matrices, at the cost of slightly worse convergence. For runtime, FP64 takes the longest, while FP32 and FP32-BF16 take a comparable amount of time. Compared to Ceres-CGNR, Graphite takes less time since it accelerates constraint-specific calculations as well. Although MegBA achieves slightly lower MSE in a few instances, Graphite uses $\frac{1}{4}$ the memory for large problems, since it avoids storing Hessian matrices, and receives a greater speedup when using single precision. These results highlight the runtime and memory advantages of performing the entire optimization on the GPU and directly utilizing the graph structure to avoid constructing intermediate matrices.

### v2

@@ 16 Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
Next, the user defines static properties for their constraints, including which vertex descriptors are involved, the data types (classes) of the observations and any non-optimizable data, the loss function used, and the differentiation mode (discussed in Section IV-B and Section IV-C). These properties also specify how to compute the residual for a constraint, and how to compute the Jacobians when using analytic differentiation (Section IV-B). 2 shows this for the point optimization problem. The user specifies a residual of size 1 (Line 3), specifies the vertex descriptor and data types (Lines 4-6), chooses the default loss function (Line 7), and chooses to use automatic differentiation (Line 8), so no Jacobian evaluation function is specified. Additionally, Lines 10-16 tell the descriptor how to compute the residual based on the equation of a circle.
1// Traits for the circle constraint
2template <typename T, typename S> struct CircleTraits {
3 static constexpr size_t dimension = 1;
4 using VertexDescriptors = std::tuple<PointDescriptor<T, S>>;
5 using Observation = T;
6 using Data = Empty; // unused - not passed to error function
7 using Loss = DefaultLoss<T, dimension>;
8 using Differentiation = DifferentiationMode::Auto;
9
10 template <typename D>
11 d_fn static void error(const D *point, const T &obs, D *error) {
12 const auto x = point[0]; // Note: A const Point<T>& can also be passed in.
13 const auto y = point[1]; // The framework automatically determines
14 const auto r = obs; // how to call the user’s function.
15 error[0] = x * x + y * y - r * r;
16 }
17};
18
19template <typename T, typename S>
20using CircleDescriptor = FactorDescriptor<T, S, CircleTraits<T, S>>;
Listing 2: Properties for the circle equation constraint descriptor.
@@
We specify properties of descriptors in this way because it lets users declare the behaviour of their types and automatically generate optimized GPU code, without relying on runtime lookups (as observed in some CPU-based optimizers [18]) to decide how each item in the graph is processed. Following Figure 1, after defining these properties, the user creates descriptors in order to construct an optimizable graph.
The user first creates a vertex descriptor for the collection of points, as shown in 3. Before doing so, they declare a graph with the desired precisions in Lines 1-3. Then in Lines 4-6, they create the vertex descriptor, reserve memory for the desired number of vertices, and add the descriptor to the graph. Then from Lines 8-10, they add a pointer for each vertex to the descriptor, and associate it with an identifier. In creating this descriptor, the user defines a batch of vertices which is processed efficiently across many GPU threads.
1using FP = float; // vertex, constraint precision
2using SP = __nv_bfloat16; // linear system precision
3Graph<FP, SP> graph; // Declare graph
4auto point_desc = PointDescriptor<FP, SP>(); // Create descriptor
5point_desc.reserve(num_vertices); // Reserve memory for num_vertices
6graph.add_descriptor(&point_desc); // Add to graph
7
8for (size_t vertex_id = 0; vertex_id < num_vertices; ++vertex_id) {
9 point_desc.add_vertex(vertex_id, &points[vertex_id]); // Add points
10}
Listing 3: Creating a vertex descriptor for 2D points.
Next, they create a descriptor for the constraints in 4. As in the previous example, they initialize the descriptor, reserve memory for the desired number of constraints (one constraint for each vertex), and add the descriptor to the graph (Lines 1-3). They initialize the default loss function on Line 4, and then create the desired number of constraints (Lines 6-8). When creating each constraint, the user specifies the identifiers of the vertices involved, the observation (the radius of the circle), the precision matrix (nullptr defaults to the identity matrix), optional constraint data (unused), and the loss function parameters. Again, by constructing this descriptor, the user enables each constraint to be processed concurrently and efficiently on the GPU.
1auto circle_desc = CircleDescriptor<FP, SP>(&point_desc); // Create descriptor
2circle_desc.reserve(num_vertices); // Reserve memory for constraints
3graph.add_descriptor(&circle_desc); // Add descriptor to graph
4const auto loss = DefaultLoss<FP, 1>(); // loss function
5
6for (size_t vertex_id = 0; vertex_id < num_vertices; ++vertex_id) {
7 circle_desc.add_factor({vertex_id}, radius, nullptr, Empty(), loss);
8}
Listing 4: Creating a descriptor for constraints.
@@
After the user finishes constructing the graph, they pass it to an optimization algorithm, along with various options, including the desired linear solver (discussed in Section IV-D), optimizing the points in-place (Figure 1). For the point optimization example, this is shown in 5. First, the user decides to fix the value of the last point they created (Line 1) and decides to disable the constraint which was created for the third point by changing its level to 1, as in g2[18] (Line 2). Next, the user configures an iterative linear solver, by selecting an identity matrix for the preconditioner (Line 3) and then instantiating the solver (Line 4) with their desired parameters. They create one stream (Line 5), since each constraint only has one vertex. Lastly, they configure the Levenberg-Marquardt algorithm on Lines 7-12, and run the optimization on Line 13.
1point_desc.set_fixed(num_vertices - 1, true); // Set the last vertex as fixed
2circle_desc.set_active(2, 0x1); // Disable third constraint for point 2
3IdentityPreconditioner<FP, SP> preconditioner;
4PCGSolver<FP, SP> solver(50, 1e-6, 10.0, &preconditioner);
5StreamPool streams(1); // chose 1 stream for 1 vertex per constraint
6
7optimizer::LevenbergMarquardtOptions<FP, SP> options;
8options.solver = &solver;
9options.initial_damping = 1e-6;
10options.iterations = 10;
11options.optimization_level = 0;
12options.streams = &streams;
13optimizer::levenberg_marquardt<FP, SP>(&graph, &options);
Listing 5: Example of optimizing noisy 2D points.
IV-B Analytic Differentiation
@@
A typical approach on the CPU is to compute each matrix for a constraint sequentially [18, 1]. However, to take advantage of the many GPU threads available, Graphite attempts to compute each Jacobian concurrently in separate streams (described in Section II). Additionally, Graphite allows these Jacobians to be evaluated dynamically [33], at the point of computation (e.g. matrix-vector multiplication), to enable matrix-free solving methods which use less memory.
IV-C Automatic Differentiation
Another method to compute Jacobians is through forward mode automatic differentiation. For example, Ceres Solver computes the gradient of a residual function using dual numbers in the form $a+b\varepsilon$, where a value is represented by real component $a$, and the partial derivative is represented by infinitesimal component $b$, where $\varepsilon^{2}=0$ [1]. To extend this for multiple inputs, Ceres implements a data structure called a Jet, which replaces the infinitesimal component with a vector [1]. Ren et. al [26] extend this concept to the GPU by proposing JetVectors, which adopt a structure of arrays data layout to enable coalesced memory transactions for better performance. However, this design is intended for optimization across multiple devices and may require more GPU memory than available on a single device [26].
@@ IV-C Automatic Differentiation
IV-D Mixed-Precision Solver
To enable mixed-precision solving, which allows the optimization to use less memory, Graphite supports the customization of floating point precisions for vertices and the linear system (Eq. 6). Essentially, these two precisions can be customized according to trade-offs between convergence, runtime performance, and memory usage. A typical approach would be to use a higher precision (double or single precision) for the vertices to maintain numerical stability, and an equal or lower precision (single precision or brain floating point) for the linear system to reduce the amount of memory needed to store it [24]. In contrast, existing solvers support only one precision at a time [26], or only allow casting the linear system to single precision [1]. As shown in Section V, using lower precisions greatly reduces the overall optimization time and the amount of GPU memory required.
As described in Section II, each iteration of Levenberg-Marquardt optimization computes a parameter step $\Delta x$. To compute this quickly on the GPU, we apply preconditioned conjugate gradients (PCG) on the Hessian, with implicit evaluation using the Jacobian matrices [33], since this does not require storing the Hessian, and comprises of algebraic operations which are largely amenable to parallelization. However, directly using PCG with lower precision types can result in poor convergence. To mitigate this, we adopt techniques which are known to help with solving larger systems. These include clamping values along the Hessian diagonal [1], rescaling the columns of each Jacobian block according to the Hessian diagonal [1, 24], and normalizing the PCG residual in each iteration [12].
In applying PCG, to improve convergence, we also use a block Jacobi preconditioner [2], because it is relatively simple to compute and does not require much memory due to its sparse block-diagonal matrix structure. We have observed that storing this preconditioner at too low of a precision can also result in convergence issues. Thus, when a 16-bit precision type such as bfloat16 is specified for the linear system, we compute and store this preconditioner at single or double precision, according to the precision of the variables.
Additionally, since the PCG solver does not require storing the Hessian matrix, we avoid explicitly constructing sparse matrices for the optimization problem. Instead, when performing an operation such as matrix-vector multiplication, the kernel uses the structure of the graph itself. This eliminates additional memory and runtime costs associated with converting and storing matrices in different formats, although each kernel must now be aware of the graph structure and whether a variable is fixed or a constraint is inactive.
@@ V Evaluation
[b]
TABLE I: Bundle adjustment results for BAL (desktop), including mean squared error, time (seconds), and GPU memory (MiB).
BAL Problem Evaluation Graphite Graphite Graphite Ceres MegBA MegBA DeepLM
Problem Size1 Metric FP64 FP32 FP32-BF16 CGNR FP64 FP32 PCG
16 Img MSE 0.43 0.43 0.43 0.43 0.43 0.43 0.43
Dubrovnik-16 22106 Pts Time 0.22 0.11 0.12 1.64 0.17 0.16 3.88
83718 Obs GPU Memory 298 282 278 336 406 352 814
21 Img MSE 1.67 1.67 1.68 1.67 1.67 1.67 1.67
Trafalgar-21 11315 Pts Time 0.11 0.069 0.068 0.82 0.098 0.088 1.61
36455 Obs GPU Memory 280 274 270 318 326 298 764
49 Img MSE 0.85 0.85 0.85 0.84 0.84 0.84 0.84
Ladybug-49 7776 Pts Time 0.063 0.041 0.046 0.66 0.086 0.074 1.14
31843 Obs GPU Memory 278 274 270 316 326 300 740
1778 Img MSE 0.67 0.67 0.67 0.67 0.66 0.67 0.67
Venice-1778 993923 Pts Time 6.00 2.80 2.87 43.71 6.96 6.15 41.67
5001946 Obs GPU Memory 1944 1190 950 2160 7810 4528 5708
4585 Img MSE 1.13 1.13 1.14 1.14 1.12 1.13 1.13
Final-4585 1324582 Pts Time 9.62 4.69 3.75 73.49 13.19 9.82 43.11
9125125 Obs GPU Memory 3102 1822 1386 3546 13882 7960 7158
13682 Img MSE 1.50 1.50 1.50 1.51 OOM OOM 1.51
Final-13682 4456117 Pts Time 26.60 11.65 11.67 210.64 OOM OOM 55.62
28987644 Obs GPU Memory 9298 5212 3828 10640 OOM OOM 15044
  • 1
    Shows the number of camera images (Img), points (Pts), and observations (Obs).
TABLE II: Bundle adjustment across differentiation modes, including mean squared error, time (seconds), and GPU memory (MiB).
@@
Final-13682 Time 26.60 417 38.81
FP64 GPU Memory 9298 3988 9298
@@
Final-13682 Time 11.65 25.22 11.46
FP32 GPU Memory 5212 2556 5212
@@
Final-13682 Time 11.67 24.48 11.27
FP32-BF16 GPU Memory 3828 2500 3828
[b]
TABLE III: Full-inertial bundle adjustment on ORB-SLAM3 maps, including the total $\chi^{2}$ error, time (seconds) and GPU memory (MiB).
TUM-VI Initial $\chi^{2}$ Graph Evaluation g2o g2o Graphite Graphite Graphite
Sequence Error Size1 Metric2 LDLT PCG FP64 FP32 FP32-BF16
@@
Orin Nano 174070 Co GPU Mem. 4674, 282 4473, 81 4449, 57
  • 1
    Shows the number of keyframes (KF), map points (MP), and constraints (Co). Each keyframe generates four variables (pose, velocity, IMU biases).
  • 2
    For the Orin Nano, we report the peak GPU memory usage and peak usage minus the baseline usage before Graphite runs.
We evaluate Graphite in terms of the error reduction, runtime performance, and GPU memory usage. First, we compare the solver in different precision modes, using small to large problems from Bundle Adjustment in the Large (BAL) [2], which serves as a common benchmark across optimization libraries. Next, we compare analytic, dynamic, and automatic differentiation modes. Lastly, we run full-inertial bundle adjustment with Graphite inside ORB-SLAM3, using maps generated from outdoors sequences in TUM-VI [28], a well-known dataset for evaluating visual-inertial SLAM systems. To assess Graphite in high-performance and resource-constrained environments, we run these experiments on a desktop machine with a 12-core Intel Core i7-12700K CPU at 3.8-4.9 GHz, 10496-core NVIDIA RTX 3090 GPU at 1.9 GHz, and 64 GB RAM, and also on an NVIDIA Jetson Orin Nano with a 6-core ARM Cortex-A78AE CPU at 1.7 GHz, 1024-core NVIDIA Ampere GPU at 1.0 GHz, and 8 GB RAM. We use unified memory to store optimizable variables.
@@ V-A Mixed-Precision Modes
We evaluate Graphite on BAL datasets using double precision (FP64), single precision (FP32), and single precision mixed with bfloat16 (FP32-BF16) on the desktop machine. We use bfloat16 rather than 16-bit half precision, because it can approximately represent the same range as single precision. For comparison, we select two general GPU-accelerated optimization libraries with BAL support and similar conjugate gradient solvers, Ceres Solver [1] and DeepLM [15], as well as MegBA [26], a fully-accelerated, high-performance bundle adjuster with single and double precision modes. Where available, we use analytic Jacobians, which have the least runtime overhead. We generate Jacobian expressions for Graphite using wrenfold [5]. To allow each optimizer sufficient opportunity to converge, we use a maximum of 50 iterations. For PCG in MegBA, the maximum number of iterations is set to 100, while 10 is used for the other configurations, since they solve the full Hessian system rather than a reduced one, which involves more work per iteration. As shown in Table I for single run experiments, Graphite uses substantially less memory when using BF16 to store the Jacobian matrices, at the cost of slightly worse convergence. For runtime, FP64 takes the longest, while FP32 and FP32-BF16 take a comparable amount of time. Compared to similar libraries, Graphite takes less time since it accelerates constraint-specific calculations as well. Although MegBA achieves slightly lower MSE in a few instances, Graphite uses $\frac{1}{4}$ the memory for large problems, since it avoids storing Hessian matrices, and receives a greater speedup when using single precision. These results highlight the runtime and memory advantages of performing the entire optimization on the GPU and directly utilizing the graph structure to avoid constructing intermediate matrices.